40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 210 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  6406   Tue Mar 13 16:56:19 2012 kiwamuUpdateLSCevolution of the sensing matrix in PRMI as a function of time

Next steps:

  • Compare the obtained sensing matrix with an Optickle model. Particularly I am interested in the absolute strengths in watts/meter
  • Noise estimation of the REFL33_Q as a MICH sensor to see if this sensor is usable for holding MICH.

Quote from #6403

Tonight I measured the sensing matrix again but this time I recorded them as a function of time using the realtime LOCKINs in the LSC front end.

The attached plots are the responses (optical gains) of PRCL and MICH in watts / meter at various sensors in time series.

  6405   Tue Mar 13 16:40:06 2012 kiwamuUpdateLSCevolution of the sensing matrix in PRMI as a function of time: details

Here I describe the measurement of the sensing matrix.

 

Motivations

  There were two reasons why I have been measuring the sensing matrix :

  1.  I wanted to know how much each element in the sensing matrix drifted as a function of time because the sensing matrix didn't agree with what Optickle predicted (#6283).
  2.  I needed to estimate the MICH responses in the 3f demodulated signals, so that I can decide which 3f signal I should use for holding MICH.

 I will report #2 later because it needs another careful noise estimation.

 

Measurement

 In order to measure the sensing matrix, the basic steps are something like this:

  1. Excite one of the DOF at a certain frequency, where a notch filter is applied in the LSC servos so that the servos won't suppress the excitation signal.
  2. Demodulate the LSC signals (e.g. C1:LSC-REFL11_I_ERR and etc.,) by the realtime LOCKINs (#6152) at the same frequency.
  3. Calibrate the obtained LOCKIN outputs to watts/meter.
In the actual measurement I choose the frequency of the excitation signal to be at 283.1 Hz,
at which any of the LSC servos don't have gains of more than 1 and there were no particular structures in the spectra.
For the amplitude of the excitation, I usually choose it to be 1000 - 2000 counts.
Because all the actuators have response functions of approximately 10-9 / f^2 meter/counts  (#5637), the actual displacement in the excited DOF should be about 10 pm level.
Therefore the excited displacements must be always in the linear ranges and also the amplitude in counts is reasonably smaller than the DAC range.
 

LOCKIN detection

The attached cartoon below shows how the LOCKIN system works for the MICH response measurement.
In the case of the PRCL response measurement, the setup is the same except that only PRM is shaken.
Here is some notes about the LOCKIN detection.
  • The LOCKIN oscillator excites ITMs differentially
    • In order to purely excites the MICH DOF, the actuation coefficients were precisely adjusted (#6398).
    • Currently ITMY has a gain of 1, and ITMX has a gain of -0.992 for the pure MICH excitation. Those numbers were put in the output matrix of the LOCKIN oscillator.
  • The demodulation phase of the LOCKIN system was adjusted to be -22 deg at the digital phase rotator.
    • This number maximizes the in-phase signals while the quadrature-phase signals give almost zero.
    • This number was adjusted when the simple MICH configuration was applied.
  • In the demodulations, the LO signals have amplitude of 100 counts to just make the demodulated signals readable numbers.

 

lockins_MICH.png

 

Calibration of the LOCKINs

  The calibration of the LOCKIN detectors is easy because all the processes takes place in the digital land, where we know all the parameters.
In this phase the goal is to calibrate the signals into counts / meter.
To calibrate the LOCKIN output signals, the following equation is used :
 
 [The obtained LOCKIN output in counts ] = H x ADOF x CLO x CEXC x 1/2  ,
 
 where H is the response of a sensor (e.g. AS55_I, AS55_Q and so on) against a particular DOF in unit of counts / m and this the quantity which we want to measure here,
ADOF is the actuator efficiency of the DOF at the excitation frequency in unit of m/counts,
CLO is the amplitude of the local oscillator signal for demodulating the sensor signals in unit of counts,
CEXC is the amplitude of the excitation signal in unit of counts,
the last 1/2 term comes from the fact there is a low pass filter in each demodulation path. 
Therefore once we measure the response of a sensor, dividing the obtained LOCKIN output by ADOF x CLO x CEXC x 1/2 gives the calibrated response in unit of counts/meter.
  ADOF are well known as they have been measured several times (#5637).
For the MICH actuator I assumed that AMICH = 2 x (ITMY response) since they are balanced through the actuation coefficients.
Note that a confirmation of this calibration has been done
when the configuration is in the simple Michelson, where we can easily estimate the response of a sensor by letting the MICH freely swing.
 

Calibration of the responses to watts/meter

  With the calibration process described above, we obtain the sensor responses in unit of counts/m.
 Then we need to do another calibration to make them into unit of W/m.
If we think about how the RFPD signal flows, we get the following gain chain.
 
[raw response in counts/m ] = Hopt x CADC x Ldemod x GWF x Ztrans x RPD
 
Hopt  is the optical gain at a sensor which we want to calibrate. It is in unit of W/m.
CADC  is the conversion factor of the ADCs and the value is CADC = 1638.4 counts/m because their resolution is 16 bit and the range is +/-20 V.
Ldemod is the conversion efficiency of the demodulation boards in unit of V/V. I used the values which Suresh measured yesterday (#6402).
GWF is the gain of the whitening filter in unit of V/V,
Ztrans is the transimpedance gain of an RFPD in unit of V/A and I used the values summarized in (the wiki),
and RPD is the responsivity of the photo diodes and I assumed RPD = 0.75 A/W for all the RFPDs.
 
Therefore the calibration can be done by dividing the raw response value by the entire gain chain of CADC x Ldemod x GWF x Ztrans x RPD.
 

Settings and parameters

  •  LSC RF demodulation phases
    •  AS55 = 17.05 deg (minimizing the PRCL sensitivity in the Q-phase)
    •  REFL11 = -41.05 deg (maximizing the PRCL sensitivity in the I-phase)
    • REFL33 = -25.85 deg (maximizing the PRCL sensitivity in the I-phase)
    • REFL55 = 4 deg (maximizing the PRCL sensitivity in the I-phase)
    • REFL165 = 39 deg (random number)
  •  Whitening filters
    • AS55 = 30 dB
    • REFL11 = 0 dB
    • REFL33 = 42 dB
    • REFL55 = 30 dB
    • REFL165 = 45 dB
  • MICH servo
    • AS55_Q for the sensor
    • G = -5 in the digital gain
    • FM2, FM3, FM5 and FM9 actiavted
    • UGF ~ 100 Hz
    • Feedback to ITMs differentially
  • PRCL servo
    • REFL33_I for the sensor
    • G = 1 in the digital gain
    • FM2, FM3, FM4, FM5 and FM9 activated
    • UGF ~ 100 Hz
    • Feedback to PRM

Quote from #6403

Tonight I measured the sensing matrix again but this time I recorded them as a function of time using the realtime LOCKINs in the LSC front end.

I will explain some more details about how I measured and calibrated the data in another elog entry.

  6404   Tue Mar 13 13:28:31 2012 Ryan FisherUpdateCDSDAQ restart with new ini file
Extra note: This was the ini file that was edited:
/cvs/cds/rtcds/caltech/c1/chans/daq/C1PEM.ini
  6403   Tue Mar 13 07:04:55 2012 kiwamuUpdateLSCevolution of the sensing matrix in PRMI as a function of time

The punch line is -- the sensing matrix still looks strange in the PRMI configuration.

 

I have been measuring the sensing matrix of the PRMI configuration because it didn't make sense (#6283).

One strange thing I have noticed before was that all the I-phase signals showed a weird behavior -- they fluctuate too much in time series.

Tonight I measured the sensing matrix again but this time I recorded them as a function of time using the realtime LOCKINs in the LSC front end.

The attached plots are the responses (optical gains) of PRCL and MICH in watts / meter at various sensors in time series.

I will explain some more details about how I measured and calibrated the data in another elog entry.

 

PRCL.png

 MICH.png

 

  6402   Mon Mar 12 22:14:56 2012 SureshUpdateRF SystemCalibration of Demod Board Efficiency.

I have completed the calibration of the demod board efficiencies.  Here is the schematic of the set-up.

 Calibration_Schematic.png

The data is given below and the data-file is attached in several different formats.

 Demod_calib.png

 

Attachment 3: Demod_calib.txt
								
	Measurements			 After corrections			Efficiency= out/in	
Demod Board	mV_ampl	mV_pk-pk	mV_pk-pk	mV_ampl	mV_ampl	mV_ampl	Vout/Vin	Vout/Vin
	PD in	Q out	I out	PD in	Q out	I out	Q out	I out
REFL33	10.6	10.0	10.0	9.4	5.0	5.0	0.53	0.53
AS11	24.0	10.0	11.0	21.3	5.0	5.5	0.23	0.26
REFL11	22.5	240.0	255.0	20.0	120.0	127.5	6.00	6.38
POX11	24.0	9.2	8.5	21.3	4.6	4.3	0.22	0.20
POY11	22.4	10.5	9.0	19.9	5.3	4.5	0.26	0.23
AS55	17.6	268.0	268.0	15.6	134.0	134.0	8.57	8.57
REFL55	19.7	15.8	15.5	17.5	7.9	7.8	0.45	0.44
POP55	18.8	278.0	274.0	16.7	139.0	137.0	8.32	8.20
REFL165	21.2	16.0	16.4	18.8	8.0	8.2	0.42	0.44
POY110	23.4	14.7	14.4	20.8	7.4	7.2	0.35	0.35
POY22	17.5	11.9	9.3	15.6	6.0	4.6	0.38	0.30
Attachment 4: Demod_calib.xlsx
  6401   Mon Mar 12 18:57:58 2012 keikoUpdateLSCRAM simulation update, RAM LSC matrix

Quote:

 I calculated the DRMI RAM LSC matrix with RAM and the operation point offsets.

  • configuration: C1 DRMI
  • RAM is added by an Mach-Zehnder ifo placed before the PRM
  • demodulation phases are optimised for each DoF
  • the operation points offset from the PDH signals are calculated and added to the optical configuration as mirror position offsets
  • Then the matrix is calculated with the offsets and the RAM
  • The set of the scrips are found as RAMmatrix.m, normMAT.m, newGetMAT.m,  on CVS/ifomodeling/40m/fullIFO_Optickle. They are a bit messy scripts at this moment.

Results:

(1) No RAM LSC matrix

  PRCL MICH SRCL
REFL11I 1 -0.001806 -0.000147
AS 55Q 0.000818 1 0.000474
AS 55 I 1.064561 902.292816 1

(2) With 1% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 1.000618 -0.001837 -0.000163
AS 55Q 0.000919 1.000521 0.000495
AS 55 I 1.169741 924.675187 1.018479
 

(3) With 5% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 0.999986 -0.001812 -0.000150
AS 55Q 0.000838 1.000028 0.000479
AS 55 I 1.084598 906.83668 1.003759
 

Adding some more results with more realistic RAM level assumption.

(4) With 0.1% RAM mod index of PM (normalized by (1) )

  PRCL MICH SRCL
REFL11I 0.99999 -0.001807 -0.000148
AS 55Q 0.000822 1.000002 0.000475
AS 55 I 1.068342 906.968167 1.00559
 

(5) With 0.5% RAM mod index of  PM (normalized by (1) )

  PRCL MICH SRCL
REFL11I  0.999978  -0.001810    -0.000149 
AS 55Q 0.000830  1.000010  0.000476 
AS 55 I 1.075926 904.321433  1.001677
 

  6400   Mon Mar 12 01:04:18 2012 keikoUpdateLSCRAM simulation update, RAM LSC matrix

 I calculated the DRMI RAM LSC matrix with RAM and the operation point offsets.

  • configuration: C1 DRMI
  • RAM is added by an Mach-Zehnder ifo placed before the PRM
  • demodulation phases are optimised for each DoF
  • the operation points offset from the PDH signals are calculated and added to the optical configuration as mirror position offsets
  • Then the matrix is calculated with the offsets and the RAM
  • The set of the scrips are found as RAMmatrix.m, normMAT.m, newGetMAT.m,  on CVS/ifomodeling/40m/fullIFO_Optickle. They are a bit messy scripts at this moment.

Results:

(1) No RAM LSC matrix

  PRCL MICH SRCL
REFL11I 1 -0.001806 -0.000147
AS 55Q 0.000818 1 0.000474
AS 55 I 1.064561 902.292816 1

(2) With 1% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 1.000618 -0.001837 -0.000163
AS 55Q 0.000919 1.000521 0.000495
AS 55 I 1.169741 924.675187 1.018479
 

(3) With 5% RAM mod index of PM (normalised by (1) )

  PRCL MICH SRCL
REFL11I 0.999986 -0.001812 -0.000150
AS 55Q 0.000838 1.000028 0.000479
AS 55 I 1.084598 906.83668 1.003759
 

  6399   Sat Mar 10 15:29:47 2012 ZachHowToComputer Scripts / ProgramsModeMatchr

For your mode matching pleasure, I have added a tool called "ModeMatchr" to the SVN under /trunk/zach/tools/modematchr/

It uses the usual fminsearch approach, but tolerates a fully astigmatic input (i.e., w0ix ≠ w0iy, z0ix ≠ z0iy) and allows for transforming to an elliptical waist  (i.e., w0fx ≠ w0fy, but z0fx = z0fy). It would be straightforward to allow for z0fx ≠ z0fy, but I have never seen a case when we actually wanted this. On the other hand, the elliptical output ability is nice for coupling to wide-angle ring cavities.

It also does the looping through available lenses for you , and retains the best solution for each lens combination in an output cell, which can then be combed with another function (getOtherSol). fminsearch is incredibly fast: with a 10-lens bank, it finds all 100 best solutions on my crappy MacBook in <10s.

I have also included the functionality to constrain the length of the total MMT to within some percentage of the optimal distance, which helps to sift through the muck .

MMrLogo.png

  6398   Sat Mar 10 02:00:03 2012 keikoUpdateLSCupdate on the locking activity

ITMX and ITMY balance for the MICH excitation (lockin) is adjusted again. Now it's ITMx = -0.992, ITMy = 1 for MICH (lockin output matrix values).

RA: what were the old values? Does this change make any difference for the signal mixing noticed before?

  6397   Fri Mar 9 20:44:24 2012 Jim LoughUpdateCDSDAQ restart with new ini file

DAQ reload/restart was performed at about 1315 PST today. The previous ini file was backed up as c1pem20120309.ini in the /chans/daq/working_backups/ directory.

I set the following to record:

The two JIMS channels at 2048:
[C1:PEM-JIMS_CH1_DQ] Persistent version of JIMS channel. When bit drops to zero indicating something bad (BLRMS threshold exceeded) happens the bit stays at zero  for >= the value of the persist EPICS variable.
[C1:PEM-JIMS_CH2_DQ] Non-persistent version of JIMS channel.

And all of the BLRMS channels at 256:
Names are of the form:
[C1:PEM-RMS_ACC1_F0p1_0p3_DQ]
[C1:PEM-RMS_ACC1_F0p3_1_DQ]

On monday I intend to look at the weekend seismic data to establish thresholds on the JIMS channels.

256 was the lowest rate possible according to the RCG manual. The JIMS channels are recorded at 2048 because I couldn't figure out how to disable the decimation filter. I will look into this further.

  6396   Fri Mar 9 16:28:10 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:
I created a page on the wiki for the new EPICS log (conlog):
https://wiki-40m.ligo.caltech.edu/aLIGO%20EPICs%20log%20%28conlog%29

I also edited this with restart instructions:
https://wiki-40m.ligo.caltech.edu/Computer_Restart_Procedures#megatron
  6395   Fri Mar 9 16:00:46 2012 steveUpdateGreen Lockinglaser emergency shut down switch replaced at the south end

Over-sized local laser emergency switch was held by large C clamp at the south end. This was replaced by a smaller one and it is mounted with magnets.

The Innolight laser was turned off, while the interlock was wired.

  6394   Fri Mar 9 15:48:56 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:

I decided to make a backup of the database and then delete it and make a new database:
 
cd ~/ryan/database_dumpMar92012
mysqldump -u root -p C1_conlog > C1_conlog.dump.Mar92012 
Note: it appears this failed the first time, thankfully this wasn't a production service yet... In the future, do not trust this backup method for important data!

Next, log into mysql as root, dump the database, remake it and grant privileges again.:
(This is saved in megatron:~/ryan/restore_database.txt
megatron:~/ryan>mysql -u root -p
Enter password:
Welcome to the MySQL monitor.  Commands end with ; or \g.
Your MySQL connection id is 174
Server version: 5.1.41-3ubuntu12.10 (Ubuntu)

Type 'help;' or '\h' for help. Type '\c' to clear the current input statement.

mysql> list databases;
ERROR 1064 (42000): You have an error in your SQL syntax; check the manual that corresponds to your MySQL server version for the right syntax to use near 'list databases' at line 1
mysql> list users;      ERROR 1064 (42000): You have an error in your SQL syntax; check the manual that corresponds to your MySQL server version for the right syntax to use near 'list users' at line 1
mysql> use C1_conlog
Reading table information for completion of table and column names
You can turn off this feature to get a quicker startup with -A

Database changed
mysql> list users;
ERROR 1064 (42000): You have an error in your SQL syntax; check the manual that corresponds to your MySQL server version for the right syntax to use near 'list users' at line 1
mysql> select User from mysql.user;                                             +------------------+
| User             |
+------------------+
| php              |
| C1_conlog_epics  |
| c1_conlog_epics  |
| root             |
| C1_conlog_epics  |
| c1_conlog_epics  |
| debian-sys-maint |
| root             |
| root             |
+------------------+
9 rows in set (0.00 sec)

mysql> show databases;                                                          +--------------------+
| Database           |
+--------------------+
| information_schema |
| C1_conlog          |
| mysql              |
+--------------------+
3 rows in set (0.00 sec)

mysql> drop database C1_conlog ;
Query OK, 2 rows affected (0.56 sec)

mysql> create database C1_conlog;
Query OK, 1 row affected (0.00 sec)

mysql> use C1_conlog ;
Database changed
mysql> SET SQL_MODE="NO_AUTO_VALUE_ON_ZERO";
Query OK, 0 rows affected (0.00 sec)

mysql>
mysql> CREATE TABLE `channels` (
    ->   `channel_id` mediumint(8) unsigned NOT NULL AUTO_INCREMENT,
    ->   `channel_name` varchar(60) NOT NULL,
    ->   PRIMARY KEY (`channel_id`),
    ->   UNIQUE KEY `channel_name` (`channel_name`)
    -> ) ENGINE=MyISAM  DEFAULT CHARSET=latin1;
Query OK, 0 rows affected (0.04 sec)

mysql>
mysql> CREATE TABLE `data` (
    ->   `acquire_time` decimal(26,6) NOT NULL,
    ->   `channel_id` mediumint(8) unsigned NOT NULL,
    ->   `value` varchar(40) DEFAULT NULL,
    ->   `status` tinyint(3) unsigned DEFAULT NULL,
    ->   `connected` tinyint(1) unsigned NOT NULL,
    ->   PRIMARY KEY (`channel_id`,`acquire_time`)
    -> ) ENGINE=MyISAM DEFAULT CHARSET=latin1;
Query OK, 0 rows affected (0.03 sec)

mysql> grant select, insert, update, execute on * to 'c1_conlog_epics'@'127.0.0.1';  Query OK, 0 rows affected (0.00 sec) 
mysql> grant select, insert, update, execute on * to 'C1_conlog_epics'@'127.0.0.1';   Query OK, 0 rows affected (0.00 sec) 
 mysql> grant select, insert, update, execute on * to 'c1_conlog_epics'@'localhost';  Query OK, 0 rows affected (0.00 sec) 
mysql> grant select, insert, update, execute on * to 'C1_conlog_epics'@'localhost';
Query OK, 0 rows affected (0.00 sec)

mysql> grant select on C1_conlog to 'php'@'%';
ERROR 1146 (42S02): Table 'C1_conlog.C1_conlog' doesn't exist
mysql> grant select on * to 'php'@'%';
Query OK, 0 rows affected (0.00 sec)

mysql> select * from mysql.users
    -> ;
ERROR 1146 (42S02): Table 'mysql.users' doesn't exist
mysql> select User from mysql.user;        
| C1_conlog_epics  |
| c1_conlog_epics  |
| root             |
| C1_conlog_epics  |
| c1_conlog_epics  |
| debian-sys-maint |
| root             |
| root             |
+------------------+
9 rows in set (0.00 sec)

mysql> Bye 



Next, I decided that I want to index on the acquire_time instead of the combination of channel_id and acquire_time (I think it makes a lot of sense for several query types, and especially debugging the conlog!):
mysql> create index acquire_time_index on data(acquire_time);
Query OK, 0 rows affected (0.04 sec)
Records: 0  Duplicates: 0  Warnings: 0

Next Fix:


The above worked well, but when I restarted the conlog, I had to re-execute the "remove_channels" from the medm, because initially all channels were being loaded (use_channel_names had all the channels still).
Additionally, there were a lot of channels with "*RMS*" in the name that were being recorded, and were changing relatively quickly, so I have added those to the remove_channel_names file.

I am going to: Backup the files in /ligo/caltech/data/conlog/c1
Edit use_channel_names to only have the good channels.
Dump the database again
Stop conlog.
Wipe the database again.
Remake the database again (with permissions and the new index).
Restart the conlog and hope!

The fix above seems to be in place and working. The database has the initial entries for the channels it monitors and is not growing without operators changing EPICs values.

  6393   Fri Mar 9 13:34:13 2012 keikoUpdateLSCupdate on the locking activity

We tried to measure the sensing matrix for MICH and PRCL last night. They look too much mixed as we expect... the matrix may be posted later. We suspect the IX and IY of the MICH excitation is not balanced very well, although Kiwamu adjusted that about two weeks ago, and it is mixing the dof. We'll try to balance it again, ans see the matrix. 

Keiko, Kiwamu

 

Quote:

[Keiko / Kiwamu]

 Some updates on the locking activity:

  • Started summarizing the data of the Michelson lock in a wiki page:
  • Gradually moving on to the PRMI lock
    • The lock stays for reasonably a long time (~20 min or more)
    • POP22/110 demod signals seemed just ADC noise.
    • A first noise budget is in process
      • The glitches make the noise level worse above 40 Hz or so in both the MICH and PRCL budgets.
    • Sensing matrix will be measured tomorrow
    • The data will be also summarized in a wiki page

 

  6392   Fri Mar 9 11:59:38 2012 Zweizig the ELOG MavenSummaryCDSNDS2 restart

 Hi Rana,

It looks like the channe list file has a few blank lines that the channel list reader is choking on. I removed the lines and it is working now.. I have made the error message a bit more obvious (gave the file name  and line number) and allowed it to ignore empty lines so this won't cause problems with future versions (when installed). The bottom line is nds2 is now running on mafalda.
Best regards,

John   

  6391   Fri Mar 9 11:02:56 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:

Too Many Fast Changing EPICS in New Conlog


I have been monitoring the new conlog, and it already has far too many rows.

I'm going through the list of channels to exclude in the update_channels script for the conlog that is currently running and removing them from
the monitored channels in the new conlog using the remove_channel_names file and the medm screen (we may want to just wipe out the tables
and start over after this is set properly, but for now I'm keeping them):
#-- Exclude a few uninteresting or obsolete categories
if ( $chan =~ m/^[BIJ]$/ ||
$chan =~ m/IOO-MC_PWR_IN/ ||
$chan eq "C1:PSL-FSS_SLOWDC" ||
$chan =~ m/PSL-STAT_.*_BITS/ ||
$chan =~ m/:IOO-PZTM[12]_(PIT|YAW)_BIAS$/ ||
$chan =~ m/DAQ.*_cycle/ ||
$chan =~ m/DAQ.*rtSeconds/ ||
$chan =~ m/C1:-.*/ ||
$chan =~ m/C1:SUP/ ||
$chan =~ m/C1:SP/ ||
$chan =~ m/C1:X/ ||
$chan =~ m/C1:TST/ ||
$chan =~ m/C1:RF/ ||
$chan =~ m/C1:UCT/ ||
$chan =~ m/C1:DU/ ||
$chan =~ m/C1:MCP/ ||
$chan =~ m/C1:MCS/ ||
$chan =~ m/C1:FEC/ ||
$chan =~ m/C1:PEM/ ||
$chan =~ m/C1:LSP/ ||
$chan =~ m/C1:NIO/ ||
$chan =~ m/C1:WFS/ ||
$chan =~ m/C1:ASC-WFS/ ||
$chan =~ m/C1:ASC-SP/ ||
$chan =~ m/C1:VG/ ||
$chan =~ m/C1:IOO-DOF/ ||
$chan =~ m/C1:IOO-EO/ ||
$chan =~ m/Name/ ||
$chan =~ m/DEFAULTNAME/ ||
$chan =~ m/:IOO-PZT.*OFFSET/ ||
$chan =~ m/PD_VAR$/ ||
$chan =~ m/_INMON$/ ||
$chan =~ m/_EXCMON$/ ||
$chan =~ m/_OUT16$/ ||
$chan =~ m/_OUTMON$/ ||
$chan =~ m/_OUTPUT$/ ||
$chan =~ m/_RSET$/ ||
$chan =~ m/_ALIVE$/ ||
$chan =~ m/VMon$/ ||
$chan =~ m/PDMon$/ ||
$chan =~ m/(BiasVMon|FE_PPOLL|MASTER_OVERFLOW|FSS_TIDALSET|CPU_LOAD|CDM
_STAT|State_Bits|INDCOFFSET)/ )

With these removals, only 15493 channels are being monitored now.
  6390   Fri Mar 9 10:44:57 2012 steveUpdateRF SystemOSA

Optical spectrum analyzers like the Attachment made by Coherent , Meles Griot- CVI and Spectral Product are all discontinued.

The 40m have Coherent models C240 analyzer with controller C251 Their Finesse measured in 2004: sn205408  F302,  sn205409  F396,

Jenne borrowed Jan's Meles Griot model 13SAE006, Peter King has the same model. FSR 300 MHZ, finnees 200 minimum

Attachment 1: OSA.pdf
OSA.pdf
  6389   Fri Mar 9 10:17:06 2012 steveHowToPEMdo not leave tables open

Type: How not to

Please do not leave optical tables open! You will be held responsible for creating dirty optics.

Attachment 1: P1080547.JPG
P1080547.JPG
  6388   Thu Mar 8 23:37:03 2012 kiwamuUpdateIOOdither Y arm dither script

I disabled the feedback to the PZT1 PITCH in the Y arm dithering scripts so that it won't push the beam away from the good point.

Currently one has to do a manual alignment only for the PZT PITCH but the rest of DOFs are still able to be automatically aligned with the script.

Quote from #6357

 The polarity for controlling the PZT1 PITCH seems to have flipped for some reason.

  6387   Thu Mar 8 17:18:19 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:
I have the aLIGO Conlog 'working' at http://192.168.113.209/conlog/conlog.php

The process is running inside a screen on megatron.

To start it running, you need to set your environment, and then run the startup.c1conlog script :

cd /cvs/cds/rtcds/caltech/c1/target/conlog/conlogepics

source conlog_environment.txt

./startup.c1conlog

This will leave you at an epics prompt, which means the code is running. (that's why I left it running in a screen for now).

To change the list of channels when the conlog is running, you need to edit the file (s):

/ligo/caltech/data/conlog/c1/add_channel_names
/ligo/caltech/data/conlog/c1/remove_channel_names

Then start up medm as follows:

cd ~/ryan/

medm -x -macro "IFO=C1" medm/CONLOG.adl

Then click the Add channel list button or Remove channel list button.

To change the channels before running the conlog from a blank database, you would edit:
/ligo/caltech/data/conlog/c1/use_channel_names (I believe this should be read whenever the conlog is restarted, but I'm only sure it does the first time you start conlog).




Documenting the rest of the installation:


Successful? Installation of Fresh EPICS and Extensions



Fresh Copy of EPICS 3.14.10


* We restarted (on Patrick's suggestion) with a fresh copy of EPICS 3.14.10 in:
/ligo/apps/linux-x86_64/epics
* I had to set a clean environment:
* Then I downloaded the tarball, unpacked it, and simply ran make within it, and it worked!
* Next, I followed Patrick's wiki instructions with only modifications to the configure/RELEASE files for the archive and ioc/conlog extensions.
* Then I realized that I had to rebuild conlog ioc after adding a directory:
/ligo/apps/linux-x86_64/epics/iocs/conlog/iocBoot/iocc1_conlog
* I had to copy the files from the h1 directory and modify them so that all reference to h1 now point to c1 in the new directory
* I then rebuilt the conlog ioc (I had to make sure setenv SCRIPTS was run again because I had logged out and back in, and I reset the whole environment properly again)

Rest of Install


* I was able to fairly trivially follow through the rest of the installation steps on Patrick's wiki, up to the "Design" section.
* Now, there is no obvious way to move forward (nothing is actually running I believe).

New Install Instructions:


"
You want to create the EPICS IOC boot directory by doing the following:

In the top level of the IOC (.../epics/iocs/conlog) with the appropriate
paths:

/epics/base/bin/linux-x86_64/makeBaseApp.pl -i -t ioc c1_conlog
What application should the IOC(s) boot?
The default uses the IOC's name, even if not listed above.
Application name? conlog

Then you will have to update the st.cmd it creates. You can compare the
st.cmd and st.cmd.backup files in the other directories to see what needs
to be changed.

I don't know if just copying the directory will work, it might.

You will also probably want to change the following line in
/epics/modules/archive/archiveApp/src/drvArchive.c:

queue = epicsMessageQueueCreate(100000000, sizeof(struct message_type));

and reduce 100000000 to something smaller depending on the amount of ram
available to the computer. I think sizeof(struct message_type) is
something like 112 bytes. Then recompile.

You basically put a file with the list of channels to use in the directory
path for "use channel list filepath" in the following command in st.cmd:
drvArchive_read_channel_list_filepaths <add channel list filepath>,<remove
channel list filepath>,<use channel list filepath>
I can send you the script that I currently use to generate that channel
list if you want, but it may need to be changed for your setup.

Once you start the ioc, open the medm which can be checked out from
subversion here: cds_user_apps/trunk/cds/common/medm/CONLOG.adl
with macro substitution for IFO: medm -x -macro "IFO=C1"
and click on the button for "Use channel list".
The number of monitored channels should increase to the number of channels
in the file.

-Patrick

...

The perl script and example file are attached, just redirect the output of
the perl script to a file. It scans autoBurt.req files in a particular
directory and its subdirectories for channel names that meet certain
criteria. All the file contains is a list of channel names, one on each
line. To start the IOC, go to the target directory and run
./startup.c1conlog.

"

{{rpfisher:scan_autoburt.pl.txt|scan_autoburt.pl.txt}}



{{rpfisher:use_channel_names.txt|use_channel_names.txt}}



My Notes Regarding These Instructions:


* Throughout the installation instructions, it probably should have been made clear that the ifo is lowercase: eg c1 (but in the end the installation mixed C1 and c1 in different places)
* Also throughout, one should be careful to replace lho with your site (eg caltech) wherever it appears
* After running the first perl script to generate the iocBoot/iocc1_conlog directory, the goal is to rebuild the whole conlog ioc by running make from epics/iocs/conlog, but before doing that:
* I needed to change the suggested line in the archive module to match the correct RAM size of the machine I am installing on (I actually gave it just less than half the free RAM), then:
* Remake the archive module
* Change into the ioc/conlog directory, remove the iocBoot I had made before for c1, rerun the perl script above, then run make from the ioc/conlog directory.
* Once that is done, you need to edit the file st.cmd to add lines for the reading and writing of channels, such as:
dbLoadRecords("db/conlog.db","IFO=C1")
drvArchive_mysql "C1_conlog","/data/mysql/C1_conlog_epics_user"
drvArchive_read_channel_list_filepaths "/ligo/caltech/data/conlog/c1/add_channel_names","/ligo/caltech/data/conlog/c1/remove_channel_names","/ligo/caltech/data/conlog/c1/use_channel_names"
drvArchive_write_channel_list_filepaths "/ligo/caltech/data/conlog/c1/channel_names","/ligo/caltech/data/conlog/c1/monitored_channel_names","/ligo/caltech/data/conlog/c1/unmonitored_channel_names"
* I also had to rerun this set of commands once that was all done:
controls: cd /opt/rtcds/<site>/<ifo>/target/conlog/conlogepics
controls: cp -r /ligo/apps/linux-x86_64/epics/iocs/conlog/db ./
controls: cp -r /ligo/apps/linux-x86_64/epics/iocs/conlog/dbd ./
controls: cp /ligo/apps/linux-x86_64/epics/iocs/conlog/iocBoot/ioc<ifo>_conlog/envPaths ./
controls: cp /ligo/apps/linux-x86_64/epics/iocs/conlog/iocBoot/ioc<ifo>_conlog/st.cmd ./
controls: echo ./bin/linux-x86_64/conlog st.cmd > startup.<ifo>conlog
controls: chmod a+x startup.<ifo>conlog
* This set of instructions seemed to be lacking, so I added this:

cp -r /ligo/apps/linux-x86_64/epics/iocs/conlog/bin/ ./

* Now the executable runs but doesn't work:
* Fixes needed:
* Need to use root permissions and make sure the files in /data/mysql have the right names for the users the code expects and also have the right passwords. (have to match capitalization appropriately for the <ifo> tag everywhere
* Might need to go into mysql and add a new user with the proper capitalization also
* Need to edit the ld_library_path to point to the new epics libraries (see the suggested environment below)
* Now, the code seems to work, but dumps me at an "epics> " prompt, I'm asking Patrick what to do next.
* I was impatient and loaded up the medm screen, and found out that the one channel I had picked was not readable (unmonitored)
* I ran a modified version of Patrick's perl script to search autoBert files for channels, and replaced my use_channel_names file with the output of the script
* Now, the medm screen shows lots of monitored channels, and the conlog is filling up! (can see it from phpmyadmin)
* Next step: I wanted to get the php pages working, so I edited the files inside /var/www/conlog:
megatron:~/ryan>diff -u /var/www/conlog/conlog.php /var/www/conlog/conlog.php.bck.Mar82012_1
--- /var/www/conlog/conlog.php  2012-03-08 15:31:53.152547771 -0800
+++ /var/www/conlog/conlog.php.bck.Mar82012_1   2012-03-08 15:28:23.062704171 -0800
@@ -19,7 +19,7 @@
        <form action="query.php" method="post">
                <h3><label for="database">Database:</label></h3>
                <select id="database" name="database">
-                       <option value="C1_conlog">C1</option>
+                       <option value="h2_conlog">h2</option>
                </select>
 
                <h3><label for="included_channels">Included channels:</label></h3>

megatron:~/ryan>diff -u /var/www/conlog/query.php /var/www/conlog/query.php.bck.Mar82012_1
--- /var/www/conlog/query.php   2012-03-08 15:33:45.122550303 -0800
+++ /var/www/conlog/query.php.bck.Mar82012_1    2012-03-08 15:32:31.772554679 -0800
@@ -168,8 +168,8 @@
        }
 
        $database_name = $_POST["database"];
-       if ($database_name == 'C1_conlog') {
-               $server = '192.168.113.209';
+       if ($database_name == 'h2_conlog') {
+               $server = 'cdsconlog';
        }
        else {
                die('Unknown database.');

* Finally, the mysql server was denying access from outside queries, so I fixed that:
megatron:~/ryan>diff -u /etc/mysql/my.cnf /etc/mysql/my.cnf.bck.Mar82012_1
--- /etc/mysql/my.cnf   2012-03-08 15:35:57.122548370 -0800
+++ /etc/mysql/my.cnf.bck.Mar82012_1    2012-03-08 15:35:10.652559315 -0800
@@ -49,7 +49,7 @@
 #
 # Instead of skip-networking the default is now to listen only on
 # localhost which is more compatible and is not less secure.
-#bind-address          = 127.0.0.1
+bind-address           = 127.0.0.1
 #
 # * Fine Tuning
 #
megatron:~/ryan>
* Now, I think everything is working * almost:
* It seems that when you first start the conlog up, it finds all the variables and inserts values of "Null" for everything, but after that it detects changes properly!


Conlog Environment


Need to source this to use the new environment:

megatron:~>cat ~/ryan/conlog_enviroment.txt
  6386   Thu Mar 8 04:13:12 2012 kiwamuUpdateLSCupdate on the locking activity

[Keiko / Kiwamu]

 Some updates on the locking activity:

  • Started summarizing the data of the Michelson lock in a wiki page:
  • Gradually moving on to the PRMI lock
    • The lock stays for reasonably a long time (~20 min or more)
    • POP22/110 demod signals seemed just ADC noise.
    • A first noise budget is in process
      • The glitches make the noise level worse above 40 Hz or so in both the MICH and PRCL budgets.
    • Sensing matrix will be measured tomorrow
    • The data will be also summarized in a wiki page
  6385   Thu Mar 8 00:57:48 2012 keikoUpdateLSCMICH noise budget on Mar 5, Mar 6, and old

Here is the recent two noise budgets of MICH, with the old measurement by Jenne. The most latest Mar 6 data is quite close to the old data, even better around 20-30 Hz. Probably some scattering source was improved?

Mar7MICHbudgettotal.png

  6384   Wed Mar 7 23:29:28 2012 keikoUpdateLSCREFL OSA observation

 kiwamu, keiko

 

 

REFLOSA.png

We measure the REFL OSA spectrum when (1) direct reflection from the PRM (2) CR lock at PRC (3) SB lock at PRC. When CR lock, both SBs are reflected from the PRC and when SB lock (ref line), some SB is sucked by PRM and looked lower than the other two lines.

 

  6383   Wed Mar 7 23:28:41 2012 Lab Cleanup CrewHowToEnvironmentTrue Beauty....

Or, how a lab should look at the end of every day.

Clean_40m_Electronics_Bench.jpg

Beat that, Bridge kids!

  6382   Wed Mar 7 22:04:05 2012 kiwamuUpdateLSCREFL OSA : how the signal look like

I was also wondering about the same thing, comparing with what Mirko obtained before with the same OSA ( #5519).

Quote from #6379

I'm puzzled why the 11MHz peak can be such high considering 1.7~2 times smaller the modulation depth.

 

  6381   Wed Mar 7 21:13:30 2012 ranaUpdateDAQNDS2

 I noticed that NDS2 was not running on mafalda as it should be. Instead, there were a couple of zombie MEDMs using up 99% of the CPU. I killed the zombies and have run the 'build channel list' script. When it finished, I tried to restart the nds server, but got the following error in the log file. Email has been dispatched to JZ.

mafalda:logs>less nds2-mafalda-201203072111.log

Configuring from file: nds2.conf
Allow list: ALL
terminate called after throwing an instance of 'std::runtime_error'
  what():  Insufficient arguments
  6380   Wed Mar 7 20:53:13 2012 keikoUpdateLSCMICH noise budget on 5 Mar

 

 Mar6-MICHbudget.png

This is the MICH noise budget on 6th March. 1Hz peak got a bit better as the BS sus control gain was increased.

 

  6379   Wed Mar 7 20:06:23 2012 KojiUpdateLSCREFL OSA : how the signal look like

I'm puzzled why the 11MHz peak can be such high considering 1.7~2 times smaller the modulation depth.

  6378   Wed Mar 7 19:10:06 2012 kiwamuUpdateLSCREFL OSA : how the signal look like

Just a quick report on the REFL OSA.

The attached plot below shows the raw signal from the REFL OSA which Keiko installed in this afternoon.

When the data was taken the beam on the REFL OSA was a direct reflection from PRM with the rest of the suspended mirrors misaligned.

One of the upper and lower 11 MHz sidebands is resolved (it is shown at 0.12 sec in the plot) while the other one is still covered by the carrier tail.

The 55 MHz upper and lower sidebands are well resolved (they are at 0.06 and 0.2 sec in the plot).

One of the oscilloscopes monitoring the OSA signals in the control room has a USB interface so that we can record the data into a USB flash memory and plot it like this.

OSArefl.png

Quote from #6375

 I swap an OSA at PSL and OSA at REFL. It was because the PSL-OSA had a better resolution, so we place this better one at REFL. The ND filter (ND3) which was on the way to REFL OSA was replaced by two BSs, because it was producing dirty multiple spots after transmitting.

 

  6377   Wed Mar 7 18:00:39 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:

Quote:

In order to install the necessary extensions to epics to make the aLIGO conlog work, I have edited one file in the base epics install that affects makefiles:

/cvs/cds/caltech/apps/linux64/epics/base/configure/CONFIG_COMMON

Jamie said he prefers diffs, so I regenerated the original file and did a diff against the current file:

megatron:configure>diff CONFIG_COMMON.orig.reconstructedMar72012 CONFIG_COMMON.bck.Mar72012
206,207c206,210
< USR_CPPFLAGS =
< USR_DBDFLAGS =
---
> USR_CPPFLAGS = -I $(EPICS_BASE)/include
> USR_CPPFLAGS += -I $(EPICS_BASE)/include/os/Linux/
> USR_CPPFLAGS += -I $(EPICS_BASE)/../modules/archive/lib/linux-x86_64/
> USR_DBDFLAGS = -I $(EPICS_BASE)/dbd
> USR_DBDFLAGS += -I $(EPICS_BASE)/../modules/archive/dbd

This is saved in CONFIG_COMMON.diff.Mar72012_1

 After following up with Patrick Thomas for a while trying to make the extensions to epics work within the currently installed epics, he decided that we should just start over with a fresh install of epics 3.14.10.

I am installing this in /ligo/apps/linux-x86_64/epics/base-3.14.10/

Prior to all of this, I had done a lot of installation and configuration of the packages needed to make LAMP work:

sudo apt-get install lamp-server^

sudo apt-get install phpmyadmin

I then continued to follow the instructions on Patrick's wiki:

https://awiki.ligo-wa.caltech.edu/aLIGO/Conlog#EDCU_library

I installed the c_string library into /ligo/apps/linux-x86_64/ according to his instructions.  (prior to my installs, there was no /ligo/ on this machine at all, so I made the needed parent directories with the correct permissions).

That should be everything up to installing epics (working on that now).

  6376   Wed Mar 7 17:39:40 2012 keikoUpdateLSCMICH noise budget on 5 Mar

 This is the calibrated MICH noise budget on Mar 5. There was a sharp peak at 1Hz and a blob on 3 Hz. The demod phase was adjusted for AS55Q.

Mar5-MICHbudget.png

 

Attachment 1: Mar5-MICHbudget.png
Mar5-MICHbudget.png
  6375   Wed Mar 7 16:32:09 2012 keikoUpdateLSCOSA

 I swap an OSA at PSL and OSA at REFL. It was because the PSL-OSA had a better resolution, so we place this better one at REFL. The ND filter (ND3) which was on the way to REFL OSA was replaced by two BSs, because it was producing dirty multiple spots after transmitting.

  6374   Wed Mar 7 15:56:36 2012 KojiUpdateLSC22/110MHz path for POP

The RF separator installed comprises of the Minicircuits filters cascaded as in the figure below.
This has one input and 4 output ports for 11, 22, 30-60, and 110MHz signal.
As seen in this entry #6167, we have 22 and 110MHz signals together with 11, 44, 66MHz signals.
They may be demodulated via a harmonic characteristic of the mixers. (Remeber mixers are not multipliers.)

RFseparator.png

 Of course the big concern is the impedance matching for those signals as usual.
The 2nd attachment shows measured impedance of the circuits with all of the ports terminated.
From the complex impedance, we can calculate the reflection coefficient. The 44 and 110MHz
components look correctly matched while the others seems largely reflected.
This certainly is not a nice situation, as the reflection can make the amplifier next to the PD unhappy
(although the reflected power is tiny in our case).

In our case more eminent problem is that the amplitude of the 22MHz signal can vary depending on the cable length by
factor of 10 in amplitude. (c.f. VSWR on the 2nd attachment.)

The transmission to each port was measured. The separation of the signals looks good. But the attenuation of the
targetted signals (i.e. insertion losses) are qulitatively consistent with the impedance. Again these losses are depend
on the cable length.

 

 

Attachment 2: impedance.pdf
impedance.pdf
Attachment 3: transmission.pdf
transmission.pdf
  6373   Wed Mar 7 13:59:07 2012 Ryan FisherSummaryComputer Scripts / ProgramsAlterations to base epics install for installing aLIGO conlog:

In order to install the necessary extensions to epics to make the aLIGO conlog work, I have edited one file in the base epics install that affects makefiles:

/cvs/cds/caltech/apps/linux64/epics/base/configure/CONFIG_COMMON

Jamie said he prefers diffs, so I regenerated the original file and did a diff against the current file:

megatron:configure>diff CONFIG_COMMON.orig.reconstructedMar72012 CONFIG_COMMON.bck.Mar72012
206,207c206,210
< USR_CPPFLAGS =
< USR_DBDFLAGS =
---
> USR_CPPFLAGS = -I $(EPICS_BASE)/include
> USR_CPPFLAGS += -I $(EPICS_BASE)/include/os/Linux/
> USR_CPPFLAGS += -I $(EPICS_BASE)/../modules/archive/lib/linux-x86_64/
> USR_DBDFLAGS = -I $(EPICS_BASE)/dbd
> USR_DBDFLAGS += -I $(EPICS_BASE)/../modules/archive/dbd

This is saved in CONFIG_COMMON.diff.Mar72012_1

  6372   Wed Mar 7 13:30:17 2012 JimUpdatePEMadded TPs and JIMS channels to PEM front-end model

[Jim Ryan]

The PEM model has been modified now to include a block called 'JIMS' for the JIMS(Joint Information Management System) channel processing. Additionally I added test points inside the BLRMS blocks that are there. These test points are connected to the output of the sqrt function for each band. I needed this for debugging purposes and it was something Jenny had requested.

The outputs are taken out of the RMS block and muxed, then demuxed just outside the JIMS block. I was unable to get the model to work properly with the muxed channel traveling up or down levels for this. Inside the JIMS block the information goes into blocks for the corresponding seismometer channel.

For each seismometer channel the five bands are processed by comparing to a threshold value to give a boolean with 1 being good (BLRMS below threshold) and 0 being bad (BLRMS above threshold). The boolean streams are then split into a persistent stream and a non-persistent stream. The persistent stream is processed by a new library block that I created (called persist) which holds the value at 0 for a number of time steps equal to an EPICS variable setting from the time the boolean first drops to zero. The persist allows excursions shorter than the timestep of a downsampled timeseries to be seen reliably.

The EPICS variables for the thresholds are of the form (in order of increasing frequency):

C1:PEM-JIMS_GUR1X_THRES1

C1:PEM-JIMS_GUR1X_THRES2

etc.

The EPICS variables for the persist step size are of the form:

C1:PEM-JIMS_GUR1X_PERSIST

C1:PEM-JIMS_GUR1Y_PERSIST

etc.

I have set all of the persist values to 2048 (1 sec.) for now. The threshold values are currently 200,140,300,485,340 for the GUR1X bands and 170,105,185,440,430 for the GUR1Y bands.

The values were set using ezcawrite. There is no MEDM screen for this yet.

PEM model was restarted at approx. 11:30 Mar. 7 2012 PST.

 

  6371   Wed Mar 7 11:44:29 2012 JenneUpdateGreen LockingXgreen beatnote cable made, laid

The Xgreen PD now has a cable going over to the beatbox. Once beatbox characterization is done I can re-find the beat, and we can do some stuff with the beatbox.

  6370   Wed Mar 7 11:20:23 2012 steveUpdateSUSoplev qpd offsets zeroed

All oplev qpd quadrons were zeroed by offset  in blocked dark condition.

  6369   Wed Mar 7 04:08:48 2012 kiwamuUpdateSUSBS SIDE gain was too low

The BS SIDE damping gain seemed too low. The gain had been 5 while the rest of the suspensions had gains of 90-500.

I increased the gain and set it to be 80.

 

I did the "Q of 5" test by kicking the BS SIDE motion to find the right gain value.

However there was a big cross coupling, which was most likely a coupling from the SIDE actuator to the POS motion.

Due to the cross coupling, the Q of 5 test didn't really show a nice ring down time series. I just put a gain of 80 to let the Q value sort of 5.

I think we should diagonalize the out matrices for all the suspensions at some point.

  6368   Tue Mar 6 23:37:31 2012 keikoUpdateSUSMICH noise budget - SUS check

Here are the OSEM spectrum of MICH suspensions (BS, IX, IY). Bounce and Roll modes are shown on 16 and 23 Hz. The filters for them has been checked.

Mar6sus1.pdf

Mar6sus2.pdf

Mar6sus3.pdf

keiko, kiwamu, Rana

Attachment 1: Mar6sus1
Attachment 2: Mar6sus2
Attachment 3: Mar6sus3
  6367   Tue Mar 6 23:14:38 2012 ranaUpdateSUSOptical Lever had bad filters

 We found that that bounce (16.1 Hz) and roll (23.5 Hz) modes on the ITMX were much higher than on the ITMY. After some checking, it seems that the bandstop filters for the

SUSPOS, SUSPIT, SUSYAW, and SUSSIDE loops are set to the correct frequencies. However, the OLPIT and OLYAW had not been set correctly. I have copied the SUS filters into the OL filterbanks and reloaded all the filter banks. Attached are the comparison of old, bad, OL with the SUS ones.

The same cockamamie situation was there for the BS & ITMY as well. Although we still don't have the roll mode frequencies listed in the mechanical resonances wiki, I have guessed that the ITMY roll frequency is the same as the ITMX, since they have nearly the same bounce frequency. OL filters for the BS & ITMY are now at the right frequency (probably). Keiko is on top of fixing things for the other optics.

I think this whole notching adventure was in Leo's hands several months ago, but WE forgot to point him at the OLs in addition to the SUS. I blame Kiwamu 50% for not supervising him and Koji by 45% for not supervising Kiwamu. The other 5% goes to someone else. You know who you are.

Attachment 1: Untitled.png
Untitled.png
  6366   Tue Mar 6 22:23:04 2012 KojiUpdateLSC22/110MHz path for POP

 

 As par Kiwamu's request, RF filters for POP22 and POP110 were installed. They are not really nice. We need to replace it with more fancy electronics.
More to come later.

 

  6365   Tue Mar 6 16:17:36 2012 JenneUpdateSUSSUS matrix diagonalization status

Has default inmat:

PRM, ITMX

Has fancy inmat:

BS, ITMY, SRM (but side is non-fancy), ETMX, ETMY, MC1, MC2, MC3

 

So it's likely that the MICH problems (giganto 1Hz peak) Keiko and Kiwamu were seeing last night had to do with ITMX having the non-optimized input matrix.  I'll try to figure out where the data from the last freeswing test is, and put in a fancy diagonalized matrix.

  6364   Tue Mar 6 16:06:15 2012 JenneConfigurationSUSPRM watchdog tripped

Quote:

ND filter ND3 (which is at the REFL port to the REFL OSA) is removed. Don't forget to put it back when you restore PRM!!!

 I don't know what tripped the PRM watchdog, but it was unhappy.  I manually moved the sliders on the IFO align screen away from the positions of the save file before turning on the damping, to make sure that I wouldn't be sending oodles of power to the REFL port, since the ND filter is still removed.  So PRM is damped now, but misaligned.

  6363   Tue Mar 6 15:22:02 2012 KeikoUpdateLSCRAM simulation update

Quote:

 I wrote an RAM simulation script ... it calculates the LSC signal offset and the operation point offset depending on the RAM modulation index.

Configuration : RAM is added on optC1, by the additional Mach-Zehnder ifo before the PRM.

Mar5RAM3.pngMar5RAM2.png

 Both are for PRCL sweep result. Note that REFL33I is always almost zero. Next step: Check the LSC matrix with matrix at the offset operation point.

 On the right figure, you see the non-zero operation points even when RAM mod index = 0. Apparently they come from non-zero loss of the model.  (Each mirror of 50ppm loss was assumed).

  6362   Tue Mar 6 01:35:03 2012 kiwamuUpdateLSCMICH characterization

[Keiko / Kiwamu]

 Update on the MICH characterization:

  • The OSAs weren't so great because the 11 MHz sidebands were covered by the carrier's tail
    • It seemed that the frequency resolution depended on the mode matching. We will try improving the mode matching tomorrow.
  • The noise budget looked very bad
    • There were huge peaks at 1 Hz, 3 Hz, 16.5 Hz and 23 Hz. Something is crazy in the vertex suspensions.
    • Keiko will post the calibrated noise budget.
  • The MICH response at AS55Q was measured and we will calibrate it into watts / meter.

 

  6361   Tue Mar 6 00:13:20 2012 keikoUpdateLSCASI signal offset

Screenshot-Untitled_Window.png

AS55Q and AS55I signals. AS55Q is around zero while AS55I has a large offset which is about the signal amplitude. It is likely because of the RAM?

keiko, kiwamu

 

 

 

 

  6360   Mon Mar 5 23:47:15 2012 keikoConfigurationLSCND2 at REFL OSA

ND filter ND3 (which is at the REFL port to the REFL OSA) is removed. Don't forget to put it back when you restore PRM!!!

  6359   Mon Mar 5 20:31:33 2012 KojiUpdateLSC22/110MHz path for POP

This a kind of self record...

We need an RF setup at POP to extract 22 and 110 MHz components separately.

I am planning to work on this in the daytime on Tuesday.

  6358   Mon Mar 5 18:12:00 2012 KeikoUpdateLSCRAM simulation update

 I wrote an RAM simulation script ... it calculates the LSC signal offset and the operation point offset depending on the RAM modulation index.

Configuration : RAM is added on optC1, by the additional Mach-Zehnder ifo before the PRM.

Mar5RAM3.pngMar5RAM2.png

 Both are for PRCL sweep result. Note that REFL33I is always almost zero. Next step: Check the LSC matrix with matrix at the offset operation point.

  6357   Mon Mar 5 17:07:58 2012 kiwamuUpdateIOOrealigned MC

I have slightly shifted the MC beam pointing to relax the PZT1 PITCH. As a result the TRY value went to 0.97 in a first lock trial.

However another issue arose:

   The polarity for controlling the PZT1 PITCH seems to have flipped for some reason.

Since it is still sort of controllable, I am leaving it as it is.

If I remember correctly, sliding the PZT1 pitch value to the positive side brought the beam spot upward in the AS CCD. But now it moves in the opposite way.

Also the ASS feedback looks tending to push the PZT1 pitch to the wrong direction.

I am not 100 % sure if the polarity really flipped, but this is my current conclusion.

 

 

(MC pointing)

  1. Locked the Y arm and aligned ITMY and ETMY with the ASS servos such that the beam spot on each test mass is well centered on the test mass.
    • With this process the eigen axis of the Y arm cavity is well prepared.
  2. Checked the beam positions of the prompt reflection light and cavity leakage field in the AS CCD.
    • It looked the incident beam needed to go upward in the CCD view.
  3. Offloaded the MC WFS feedback values to the MC suspension DC biases in a manual way.
  4. Disabled the MC WFS servos. The MC transmitted light didn't become worse, which means the suspensions were well aligned to the input beam
  5. Changed the DC bias in the MC2 PITCH, to bring the beam spot upward. I changed the DC bias by ~ 0.1 or in the EPICS counts.
  6. Aligned the zig-zag steering mirrors on the PSL table to match the incident beam to the new MC eigen beam axis.
    • The transmitted DC light and reflected DC values went back to 27000 counts and 0.58 counts respectively without the WFS servos.
  7. Re-engaged the WFS servos.

Quote from #6351

PZT1 started railing in the pitch direction and because of this TRY doesn't go more than 0.7. I will leave it as it is for tonight.

Tomorrow I will shift the alignment of the MC to make the PZT1 happier. 

 

ELOG V3.1.3-